# Secret extraction attacks against obfuscated IQP circuits

### v1

@@ 3.1 Hiding a secret string in an IQP circuit
    the radical is doubly even, i.e., for all $\mathbf{x}\in\operatorname{rad}\mathcal{C}_{s}$, $|\mathbf{x}|=0\mod 4$.
@@ 4.1.3 Reconstructing the support from random samples
We proceed to the third ingredient of the analysis—asking whether the support the numerically obtained elements of the radical is likely to be equal to the obfuscated first $m_{1}$ coordinates.
@@
While it is highly plausible that the uniform random model accurately captures the distribution of $\mathbf{H}(\ker\mathbf{G})$, this is less true for $\mathbf{D}$, which is constrained to have doubly-even, orthogonal and linearly independent columns. Nonetheless, it will turn out that the predictions made based on this model fit the empirical findings very well. This suggests that in the choice of $\mathbf{D}$, full support on $[m_{1}]$ is attained at least as fast as suggested by the random model. We leave finding a theoretical justification for this behavior as an open question.222 One difficulty to overcome is that it is not apparent that sequential sampling algorithms, like the one implemented in [23], produce a uniform distribution over all subspaces compatible with the geometric constraints. Usually, such results are proven by invoking a suitable version of Witt’s Lemma to establish that any partial generator matrix can be extended to a full one in the same number of ways. In geometries that take Hamming weight modulo $4$ into account, there may however be obstructions against such extensions. A relevant reference is [26, Sec. 4] (see also Refs. [27, 28] for a discussion in the context of stabilizer theory). The theorem in that section states that two isometric subspaces can be mapped onto each other by a global isometry only if they both contain the all-ones vector or if neither of them does. The resulting complications have prevented us from finding a simple rigorous version of Lemma 4 that applies to random generator matrices of doubly-even spaces.
@@ 5.1.1 Analysis
$\mathbf{H}_{s}\mathbf{d}\in\operatorname{rad}\mathcal{C}_{s}\quad\Rightarrow\quad\mathbf{s}\in\ker(\mathbf{G}_{\mathbf{d}}).$ (15)
@@
By definition, it holds that $\mathbf{H}_{s}\mathbf{d}\in\operatorname{rad}\mathcal{C}_{s}$ iff
$\mathbf{d}^{T}\mathbf{H}_{s}^{T}\mathbf{H}_{s}\mathbf{e}=\mathbf{d}^{T}\mathbf{H}^{T}\mathbf{H}_{s}\mathbf{e}=0\qquad\forall\,\mathbf{e}\in\mathbbm{F}_{2}^{n}.$ (16)
Because $\mathbf{s}\mapsto\mathbf{H}_{\mathbf{s}}$ is linear, the above means that every element $\mathbf{H}{\mathbf{d}}\in\operatorname{rad}\mathcal{C}_{s}$ of the radical gives rise to a set of linear equations (one for each $\mathbf{e}\in\mathbbm{F}_{2}^{n}$) for the secret $\mathbf{s}$. These equations can be compactly written as
$\mathbf{d}^{T}\mathbf{H}^{T}\mathbf{H}_{s}=0\quad\Leftrightarrow\quad\mathbf{H}^{T}\mathbf{H}_{\mathbf{d}}\mathbf{s}=0\quad\Leftrightarrow\quad\mathbf{H}_{\mathbf{d}}^{T}\mathbf{H}_{\mathbf{d}}\mathbf{s}=0.$ (17)
@@
(b) Dimension of $\bigcap_{i\in[k]}\ker(\mathbf{G}_{\mathbf{d}^{i}})$ (green) for the QRC construction with $q=103,n=q+r,r=(q+1)/2,m=2q$. We used 100 random instances and 1000 random choices of $\mathbf{d}^{1},\ldots,\mathbf{d}^{k}$ per point. The simple theoretical prediction of $2^{-k}(n-m/2)$ (pink) is seen to be in good agreement with the numerical experiments up to a constant prefactor.
@@ 5.2 The Double Meyer Attack
In the previous section, we discussed how to exploit statistical fluctuations to avoid having to search through large kernels. As $m-n/2$ increases, this strategy will fail to be effective. We thus present another ansatz, the Double Meyer Attack, stated as Algorithm 3. It reduces the size of the kernel, essentially by running several Linearity Attacks at once.
@@ 5.2.1 Analysis
At a high-level, the Double Meyer Attack examines the elements of the intersection of the kernels of several $\mathbf{G}_{\mathbf{d}^{i}}$, where the $\mathbf{d}^{i}$ are uniformly distributed random vectors.333 The essential generalization over Kahanamoku-Meyer’s Linearity Attack can be gleamed from the simplest special case $k=2$, which, of course, explains the name we have adopted for this ansatz. Since the events $\mathbf{H}_{\mathbf{s}}\mathbf{d}^{i}\in\operatorname{rad}\mathcal{C}_{s}$ are independent for different $i$’s, we have that
@@
At the same time, $\dim(\bigcap_{i\in[k]}\ker\mathbf{G}_{\mathbf{d}^{i}})\approx 2^{-k}(m-n/2)$, a behavior suggested by the fact that any given row of $\mathbf{H}$ will have non-zero inner product with all the $\mathbf{d}^{i}$ with probability $2^{-k}$.
@@ 5.3 Hamming’s Razor
In this section, we describe a method that allows one to “shave off” certain rows and columns from $\mathbf{H}$ without affecting the code space $\mathcal{C}_{\mathbf{s}}$. Such redundancies can be identified given a vector $\mathbf{d}\in\mathbbm{F}_{2}^{n}$ such that $\mathbf{H}\mathbf{d}$ has low Hamming weight. The method comes in two varieties: The simpler Singleton Razor, discussed first, and the more general Hamming’s Razor.
@@ 5.3.2 Hamming’s Razor
This suggests to choose $S$ by including each coordinate $i\in[m]$ with probability $p$ such that $pm_{2}>m_{2}-(n-g-d)$ and $pm_{1}<k_{\infty}$. For the challenge parameters, these requirements are compatible with the range $[0.01,0.13]$ for $p$.
@@ 6 Discussion
We emphasize that, in our judgment, the problem of finding ways to efficiently certify the operation of near-term quantum computing devices is an important one, and the idea to use obfuscated quantum circuits remains appealing. More generally, results that exhibit weaknesses in published constructions should not cause the community to turn away, but should rather serve as sign posts guiding the way to more resilient schemes. It may be instructive to compare the situation to the more mature field of classical cryptography, which benefits from a large public record of cryptographic constructions and their cryptanalysis. New protocols can thus be designed to resist known exploits and be vetted against them. The well-documented story of differential cryptanalysis provides an instructive example.
We hope that the quantum information community will look at the present results in a similarly constructive spirit.
@@ Acknowledgements
First and foremost, we are grateful to Michael Bremner, Bin Cheng, and Zhengfeng Ji for setting the challenge and providing a thorough code base that set a low threshold for its study. We feel it is worth stating that we have tried numerous times to construct alternative hiding schemes ourselves—but found each of our own methods much easier to break than theirs.
We like to thank Bin Cheng for very helpful comments on a draft version of this paper.
@@
DG’s contribution has been supported by the German Ministry for Education and Research under the QSolid project, and by Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 (390534769). DH acknowledges funding from the US Department of Defense through a QuICS Hartree fellowship.

### v2

@@ 3.1 Hiding a secret string in an IQP circuit
    the radical is doubly even, i.e., for all $\mathbf{x}\in\operatorname{rad}\mathcal{C}_{\mathbf{s}}$, $|\mathbf{x}|=0\mod 4$.
@@ 4.1.3 Reconstructing the support from random samples
We proceed to the third ingredient of the analysis—asking whether the support of the numerically obtained elements of the radical is likely to be equal to the obfuscated first $m_{1}$ coordinates.
@@
While it is highly plausible that the uniform random model accurately captures the distribution of $\mathbf{H}(\ker\mathbf{G})$, this is less obvious for $\mathbf{D}$, which is constrained to have doubly-even, orthogonal and linearly independent columns. Nonetheless, it will turn out that the predictions made based on this model fit the empirical findings very well. This suggests that in the choice of $\mathbf{D}$, full support on $[m_{1}]$ is attained at least as fast as suggested by the random model. We leave finding a theoretical justification for this behavior as an open question.222 One difficulty to overcome is that it is not apparent that sequential sampling algorithms, like the one implemented in [23], produce a uniform distribution over all subspaces compatible with the geometric constraints. Usually, such results are proven by invoking a suitable version of Witt’s Lemma to establish that any partial generator matrix can be extended to a full one in the same number of ways. In geometries that take Hamming weight modulo $4$ into account, there may however be obstructions against such extensions. A relevant reference is [26, Sec. 4] (see also Refs. [27, 28] for a discussion in the context of stabilizer theory). The theorem in that section states that two isometric subspaces can be mapped onto each other by a global isometry only if they both contain the all-ones vector or if neither of them does. The resulting complications have prevented us from finding a simple rigorous version of Lemma 4 that applies to random generator matrices of doubly-even spaces.
@@ 5.1.1 Analysis
$\mathbf{H}_{\mathbf{s}}\mathbf{d}\in\operatorname{rad}\mathcal{C}_{\mathbf{s}}\quad\Rightarrow\quad\mathbf{s}\in\ker(\mathbf{G}_{\mathbf{d}}).$ (15)
@@
By definition, it holds that $\mathbf{H}_{\mathbf{s}}\mathbf{d}\in\operatorname{rad}\mathcal{C}_{\mathbf{s}}$ iff
$\mathbf{d}^{T}\mathbf{H}_{\mathbf{s}}^{T}\mathbf{H}_{\mathbf{s}}\mathbf{e}=\mathbf{d}^{T}\mathbf{H}^{T}\mathbf{H}_{\mathbf{s}}\mathbf{e}=0\qquad\forall\,\mathbf{e}\in\mathbbm{F}_{2}^{n}.$ (16)
Because $\mathbf{s}\mapsto\mathbf{H}_{\mathbf{s}}$ is linear, the above means that every element $\mathbf{H}{\mathbf{d}}\in\operatorname{rad}\mathcal{C}_{\mathbf{s}}$ of the radical gives rise to a set of linear equations (one for each $\mathbf{e}\in\mathbbm{F}_{2}^{n}$) for the secret $\mathbf{s}$. These equations can be compactly written as
$\mathbf{d}^{T}\mathbf{H}^{T}\mathbf{H}_{\mathbf{s}}=0\quad\Leftrightarrow\quad\mathbf{H}^{T}\mathbf{H}_{\mathbf{d}}\mathbf{s}=0\quad\Leftrightarrow\quad\mathbf{H}_{\mathbf{d}}^{T}\mathbf{H}_{\mathbf{d}}\mathbf{s}=0.$ (17)
@@
(b) Dimension of $\bigcap_{i\in[k]}\ker(\mathbf{G}_{\mathbf{d}^{i}})$ (green) for the QRC construction with $q=103,n=q+r,r=(q+1)/2,m=2q$. We used 100 random instances and 1000 random choices of $\mathbf{d}^{1},\ldots,\mathbf{d}^{k}$ per point. The simple theoretical prediction of $2^{-k+1}(n-m/2)$ (dotted pink) is seen to be in good agreement with the numerical experiments up to a constant prefactor.
@@ 5.2 The Double Meyer Attack
In the previous section, we discussed how to exploit statistical fluctuations to avoid having to search through large kernels. But as $n-m/2$ increases, this strategy will eventually fail to be effective. We thus present another ansatz, the Double Meyer Attack, stated as Algorithm 3. It reduces the size of the kernel, essentially by running several Linearity Attacks at once.
@@ 5.2.1 Analysis
At a high-level, the Double Meyer examines the elements of the intersection of the kernels of several $\mathbf{G}_{\mathbf{d}^{i}}$, where the $\mathbf{d}^{i}$ are uniformly distributed random vectors 333 The essential generalization over Kahanamoku-Meyer’s Linearity Attack can be gleamed from the simplest special case $k=2$, which, of course, explains the name we have adopted for this ansatz. . Since the events $\mathbf{H}_{\mathbf{s}}\mathbf{d}^{i}\in\operatorname{rad}\mathcal{C}_{\mathbf{s}}$ are independent for different $i$’s, we have that
@@
At the same time, we expect $\dim(\bigcap_{i\in[k]}\ker\mathbf{G}_{\mathbf{d}^{i}})\approx 2^{-k+1}(n-m/2)$. To see this, observe that the kernel of $\mathbf{G}_{\mathbf{d}^{i}}$ decomposes as $\ker\mathbf{G}_{\mathbf{d}^{i}}=\ker\mathbf{H}_{\mathbf{d}^{i}}+\operatorname{rad}\operatorname{range}(\mathbf{H}_{\mathbf{d}^{i}})$. For random vectors $\mathbf{d}^{i}$ we then expect $\operatorname{rad}\operatorname{range}(\mathbf{H}_{\mathbf{d}^{i}})$ to be iid. random subspaces, while the kernels of $\mathbf{H}_{\mathbf{d}^{i}}$ are correlated, since every pair of $\ker\mathbf{H}_{\mathbf{d}^{1}},\ker\mathbf{H}_{\mathbf{d}^{2}}$ shares half the rows. Thus, the intersection of $\operatorname{rad}\operatorname{range}\mathbf{H}_{\mathbf{d}^{i}}$ decays exponentially with $k$, while the intersection $\bigcap_{i\in[k]}\ker\mathbf{H}_{\mathbf{d}^{i}}$ decreases much faster as $n-(1-2^{-k})m$. Altogether, the intersection decomposes into sums of intersections (for $k=2$)
$\bigcap_{i=1}^{2}\ker\mathbf{G}_{\mathbf{d}^{i}}=(\ker\mathbf{H}_{\mathbf{d}^{1}}+\operatorname{rad}\operatorname{range}(\mathbf{H}_{\mathbf{d}^{1}}))\cap(\ker\mathbf{H}_{\mathbf{d}^{2}}+\operatorname{rad}\operatorname{range}(\mathbf{H}_{\mathbf{d}^{2}})),$ (23)
where the exponential decay as $2^{-k}$ stemming from the last term dominates the scaling.
@@ 5.2.2 Application to the Extended QRC construction
5.2.3 Further improvements
As stated, the Double Meyer draws vectors $\mathbf{d}^{i}$ uniformly at random, in the hope that $\mathbf{H}_{\mathbf{s}}\mathbf{d}^{i}\in\operatorname{rad}\mathcal{C}_{\mathbf{s}}$. But there might be more efficient ways of obtaining vectors $\mathbf{d}^{i}$ satisfying this condition. For instance, under the assumptions of Corollary 3.1, any element $\mathbf{d}$ of $\mathbf{H}(\ker(\mathbf{H}^{T}\mathbf{H}))$ has this property. Adding such vectors to the collection of $\mathbf{d}^{i}$’s therefore provides additional constraints for $\mathbf{s}$ at essentially no computational cost.
In particular, this modification of the Double Meyer would break a variant of the construction Bremner et al. [31] proposed as an initial reaction to the pre-print of this paper with parameters --AB-type zero --concat_D --concat_C1 as implemented in commit 7d3bd3 of their github repository [32].
@@ 5.3 Hamming’s Razor
In this section, we describe a method that allows one to “shave off” certain rows and columns from $\mathbf{H}$ without affecting the code space $\mathcal{C}_{\mathbf{s}}$. Such redundancies can be identified given a vector $\mathbf{d}\in\mathbbm{F}_{2}^{n}$ such that $\mathbf{H}\mathbf{d}$ has low Hamming weight. The method comes in two varieties: The simpler Singleton Razor, discussed first, and the more general Hamming’s Razor proper.
@@ 5.3.2 Hamming’s Razor
This suggests to construct $S$ by including each coordinate $i\in[m]$ with probability $p$, chosen such that $pm_{2}>m_{2}-(n-g-d)$ and $pm_{1}<k_{\infty}$. For the challenge parameters, these requirements are compatible with the range $[0.01,0.13]$ for $p$.
@@ 6 Discussion
As a reaction to a draft version of this paper, Bremner, Cheng, and Ji have modified their proposal to evade the attacks described here. A first update (communicated privately) led to our improved Double Meyer, as sketched in Sec. 5.2.3. As of September 2024, the authors have provided us with a version of their protocol in which we have not found a weakness [33].
It may be instructive to compare the situation to the more mature field of classical cryptography, which benefits from a large public record of cryptographic constructions and their cryptanalysis. New protocols can thus be designed to resist known exploits and be vetted against them. In particular, most cryptographic protocols actually in use are not rigorously proven to be secure. Instead, trust in them is based on a long and public history of constructions, as well as successful and unsuccessful attempts at attacking them. The well-documented story of differential cryptanalysis provides an instructive example.
In this light, we consider the high-level contributions of our paper to be this: 1. It shows potential users of crytpographically backed-up demonstrations of quantum supremacy that previous proposals have been broken repeatedly, so that their security should not be taken for granted. 2. The fact that we have not yet been able to identify an efficient attack against the latest version of the protocol should raise one’s trust in that version of the proposal, compared to a protocol that has not been the subject of a public security review. 3. Our attacks clarify properties of the IQP-based protocol that make it amenable to classical attacks as well as a collection of cryptanalyltic techniques exploiting those. These tools can guide and must be taken into account by designers of future constructions.
At the same time, our cryptanalysis remains to have two important consequences on IQP-based verified quantum advantage using the construction of Bremner et al. [23]. First, our Double Meyer attack remains valid for all instances and has quasi-polynomial runtime $2^{-\Omega(\log^{2}n)}$ given the verification condition that the signal $2^{-g}$ remains inverse polynomially large. This is a significant improvement over the previous state of the art, which was an exponential-time algorithm. Second, in order to circumvent our attacks, Bremner et al. [33] have had to significantly increase the problem or ‘key’ size from the original challenge data with parameters $n=300,m=360,g=4$ to $n=700$ qubits, $m=1200$ Hamiltonian terms and a verification signal of $2^{-g}=2^{-10}$. It is therefore an interesting open problem to compare the implementation cost of this scheme to other resource-efficient schemes which come with provable security guarantees based on well-studied classical assumptions [5, 6].
We emphasize that, in our judgment, the problem of finding ways to efficiently certify the operation of near-term quantum computing devices is an important one and the idea to use obfuscated quantum circuits remains appealing. More generally, the story of IQP-based verified quantum advantage and our contribution to it illustrate that results that exhibit weaknesses in published constructions should not cause the community to turn away, but should rather serve as sign posts guiding the way to more resilient schemes. We remain curious whether the security of the new construction of Bremner et al. [23] holds up to further scrutiny by the community.
@@ Acknowledgements
First and foremost, we are grateful to Michael Bremner, Bin Cheng, and Zhengfeng Ji for setting the challenge and providing a thorough code base that set a low threshold for its study. We feel it is worth stating that we have tried numerous times to construct alternative hiding schemes ourselves—but found each of our own methods much easier to break than theirs. We also thank Mick, Bin, and Zhengfeng for detailed comments and discussions on earlier versions of this paper.
@@
DG’s contribution has been supported by the German Ministry for Education and Research under the QSolid project, and by Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 (390534769). DH acknowledges funding from the US Department of Defense through a QuICS Hartree fellowship and from the Simons Institute for the Theory of Computing, supported by DOE QSA.